At STP, between 4.0 grams of helium and 44.0 grams of carbon dioxide, what can you say about the number of molecules in each sample of gas? Explain.
1
Expert's answer
2020-04-06T11:46:50-0400

In 4.0 grams of He there is 4/2 = 2 mole of helium, i.e. 2*6.02*10^23 = 1.204*10^24 molecules

In 44.0 grams of CO2 there is 1 mole, i.e. 6.02*10^23 molecules